OpenStudy (anonymous):

f(x)=cosx+2cos^2x defined for pi/2

OpenStudy (anonymous):

\[f(x)=\cos(x)+2\cos^2(x)\] ?

OpenStudy (anonymous):

is that the function?

OpenStudy (anonymous):

yup!

OpenStudy (anonymous):

did you find the derivative?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

i got \[f'(x)=\sin(x)(1-4\cos(x))\]

OpenStudy (anonymous):

I got (-sinx)(2cosx)

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

\[-\sin(x)-4\cos(x)\sin(x)\] is a start right?

OpenStudy (anonymous):

maybe the problem is with the derivative of \(2\cos^2(x)\)

OpenStudy (anonymous):

by the chain rule it is \(-4\cos(x)\sin(x)\)

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

actually yeah

OpenStudy (anonymous):

ok then set it equal to zero and solve it will be goofy

OpenStudy (anonymous):

uh... i dont even know where to start

OpenStudy (anonymous):

\[f'(x)=\sin(x)(1-4\cos(x))\]\[\sin(x)(1-4\cos(x))=0\]

OpenStudy (anonymous):

right. I got up to there lol

OpenStudy (anonymous):

so \(\sin(x)=0\) which means in your interval \(x=\frac{\pi}{2}\) which is not in fact in your interval is it?

OpenStudy (anonymous):

so the other possibility is that \[1-4\cos(x)=0\] giving \[\cos(x)=\frac{1}{4}\]

OpenStudy (anonymous):

as for that one, \(x=\cos^{-1}(\frac{1}{4})\) and i have no idea what that is , use a calculator or just leave it at that

OpenStudy (anonymous):

oh damn i made an algebra mistake!!

OpenStudy (anonymous):

it is \[\sin(x)(-1-4\cos(x))=0\]

OpenStudy (anonymous):

so \[\cos(x)=-\frac{1}{4}\] or \[x=\cos^{-1}(-\frac{1}{4})\]

OpenStudy (anonymous):

sorry about that

OpenStudy (anonymous):

thanks!